OpenStudy (anonymous):

if I and V are vectors of current and voltage.Im asked to find I*.V. What does that means? What operation should I carry out between I and V?. please help

OpenStudy (anonymous):

You mean dot product?

OpenStudy (anonymous):

dont know. How do I carry out dot product? can you give me a formula?

OpenStudy (anonymous):

if vector a=a1+a2+a3 and b=b1+b2+b3 then dot product ab=a1*b1+a2*b2+a3*b3 In your case, those would be I1+I2+I3 and V1+V2+V3 (all these numbers are sub indices but it's a pain to draw, but you know what I mean)

OpenStudy (anonymous):

you mean, I have sto multiply I1*V1 + I2*V2 + ... InVn?

OpenStudy (anonymous):

Yes. Also, very important, you multiply only the coefficients and then add them up. The result of dot product is a scalar, NOT a vector.

OpenStudy (anonymous):

When I do that I get something really weird This operation is meant to be the power in one device. The weird part of it is that I end up with a negative value. I dont think power can be negative. Wrong?

OpenStudy (anonymous):

If I=5i+2j and V=3i-4j then dot product is: 5*3+2*(-4)=15-8=7

OpenStudy (anonymous):

Actually it can. If the device is sucking in power it's negative. If the device is providing power it's positive.

OpenStudy (anonymous):

power can be negative, it means power is being supplied instead of being used

OpenStudy (anonymous):

imranmeah91 and I disagree. So I'm gonna say I'm probably wrong. The bottom line is: yes, power can be negative. Depends on the convention of sings you're following.

OpenStudy (anonymous):

Rudy is exactly right, negative and positive does not matter; what matters in difference in signs.

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (anonymous):

If I have the eficient current and the eficient voltage. Can I found apparent power by multiplying Ie and Ve like this: Ie=318,198 Ve=155,563 P=Ie*Ve What are the units of aparent power btw?

OpenStudy (anonymous):

Volt-Ampers or KiloVolt-Amperes or (you might have read on a transformer) KVA